Session: Oct 2007 (session 3)
Wavelength: 6 cm
To give feedback on this experiment, click
Onsala (Michael Lindqvist at Sat Nov 3 15:36:21 2007) Failure :
Could only observe until around 05:20 UTC due to high wind speed (> 13 m/s).
Noto (P. Cassaro at Sun Nov 4 00:30:54 2007) Success :
No known problems
Torun (K. Borkowski at Mon Nov 5 08:42:48 2007) Minor failures :
BBC showed 'unlock' for a few brief times. No other known problems.
Effelsberg (A. Polatidis at Mon Nov 5 10:37:27 2007) Success :
Lost scans Day 307 UT 0514-0519 for pointing, 1100-1140, 1205-1208 due to Mk5 server failure. Otherwise no problems.
Medicina (Andrea Orlati at Tue Nov 6 09:09:56 2007) Minor failures :
The Pcal was on.
Jodrell_Bank (Alastair Gunn at Tue Nov 6 10:07:04 2007) Success :
The MkII telescope was used (Jb2).
No known problems.